d^2y/d^2 = 1/x(x+1) + cosec^2x
1
Expert's answer
2019-10-31T10:56:18-0400

We need to solve the equation


"y''(x) = \frac{1}{{x(x + 1)}} + \frac{1}{{{{\sin }^2}x}}"

(we used that "{\csc ^2}x = \frac{1}{{{{\sin }^2}x}}")

To solve it, we need to integrate it twice. Let's do the first integration


"y'(x) = {C_1} + \int {\frac{{dx}}{{x(x + 1)}}} + \int {\frac{{dx}}{{{{\sin }^2}x}}}"

To find the integral "\int {\frac{{dx}}{{x(x + 1)}}}" let's represent the fraction "\frac{1}{{x(x + 1)}}" as a sum of simple fractions "\frac{1}{{x(x + 1)}} = \frac{1}{x} - \frac{1}{{x + 1}}" then


"\int {\frac{{dx}}{{x(x + 1)}}} = \int {\frac{{dx}}{x}} - \int {\frac{{dx}}{{x + 1}}} = \ln \left| x \right| - \ln \left| {x + 1} \right|"

The nex integral "\int {\frac{{dx}}{{{{\sin }^2}x}}}" can be found if we you notice that


"\frac{{d\cot (x)}}{{dx}} = \frac{{ - \sin x}}{{\sin x}} - \frac{{\cos x}}{{{{\sin }^2}x}}\cos x = - 1 - \frac{{1 - {{\sin }^2}x}}{{{{\sin }^2}x}} = - \frac{1}{{{{\sin }^2}x}}"

It means that

"\int {\frac{{dx}}{{{{\sin }^2}x}}} = - \cot x"

We get

"y'(x) = {C_1} + \ln \left| x \right| - \ln \left| {x + 1} \right| - \cot x"

Do the integration again


"y(x) = {C_1}x + {C_2} + \int {\ln \left| x \right|dx} - \int {\ln \left| {x + 1} \right|dx} - \int {\cot xdx}"

First integral can be calculated by parts. Let "x > 0" we get


"\int {\ln xdx} = x\ln x - \int {x \cdot \frac{{dx}}{x}} = x(\ln x - 1)+c"

For "x < 0" using "\xi = - x" we get


"\int {\ln ( - x)dx} = - \int {\ln \xi d\xi } = - \xi (\ln \xi - 1)+c = x(\ln ( - x) - 1)+c"

Thus


"\int {\ln \left| x \right|dx} = x(\ln \left| x \right| - 1)+c"

The second integral can be reduced to the first using "\eta = x + 1"


"\int {\ln \left| {x + 1} \right|dx} = \int {\ln \left| \eta \right|d\eta }"

Thus


"\int {\ln \left| {x + 1} \right|dx} = (x + 1)(\ln \left| {x + 1} \right| - 1)+c"

The last integral


"\int {\cot xdx} = \int {\frac{{\cos x}}{{\sin x}}dx} = \int {\frac{{d(\sin x)}}{{\sin x}}} = \ln \left| {\sin x} \right|+c"

And we get (we omit some terms because they can be included in "{C_1}x" and "{C_2}")


"y(x) = {C_1}x + {C_2} + x\ln \left| x \right| - (x + 1)\ln \left| {x + 1} \right| - \ln \left| {\sin x} \right|"